Find the breadth of a rectangular plot of land, if its area is 440 m2 and the length is 22 m. Also find its perimeter

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

From the question it is given that,
Area of the rectangular plot = 440 m^2
Length of the rectangular plot = 22 m
We know that,
Area of the rectangle = Length × Breadth
440 = 22 × Breadth
Breadth = 440/22
Breadth = 20 m
Then,
Perimeter of the rectangle = 2(Length + Breadth)
= 2 (22 + 20)
= 2(42)
= 84 m
∴Perimeter of the rectangular plot is 84 m.

Thanks